### v1

# Pathways of Thoughts: Multi-Directional Thinking for Long-form Personalized Question Answering

Abstract.
Personalization is essential for adapting question answering (QA) systems to user-specific information needs, thereby improving both accuracy and user satisfaction. However, personalized QA remains relatively underexplored due to challenges such as inferring preferences from long, noisy, and implicit contexts, and generating responses that are simultaneously correct, contextually appropriate, and aligned with user expectations and background knowledge. To address these challenges, we propose Pathways of Thoughts (PoT), an inference-stage method that applies to any large language model (LLM) without requiring task-specific fine-tuning. The approach models the reasoning of an LLM as an iterative decision process, where the model dynamically selects among cognitive operations such as reasoning, revision, personalization, and clarification. This enables exploration of multiple reasoning trajectories, producing diverse candidate responses that capture different perspectives. PoT then aggregates and reweights these candidates according to inferred user preferences, yielding a final personalized response that benefits from the complementary strengths of diverse reasoning paths. Experiments on the LaMP-QA benchmark for personalized QA show that PoT consistently outperforms competitive baselines, achieving up to a 13.1% relative improvement. Human evaluation corroborates these results, with annotators preferring outputs from PoT in 66% of cases and reporting ties in only 15% of cases.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Personalizing large language models (LLMs) has emerged as an important topic due to its applications in content generation, writing assistance, education, and recommendation (Salemi et al., 2024b; Kumar et al., 2024; Mysore et al., 2023; Li et al., 2023; Naumov et al., 2019). By conditioning outputs on individual users’ preferences, contexts, and needs, personalized text generation improves user engagement, satisfaction, and overall effectiveness (Salemi et al., 2025). Personalization is especially valuable in information-seeking systems, including agentic web search and question answering–based information access (Salemi and Zamani, 2025a, b). For instance, the response to a question like “What is the best strategy to study for a difficult exam?” depends on various user-specific factors, including their background, prior knowledge, learning style, and habits. A tailored response can recommend strategies that align with the user’s cognitive strengths, subject matter, and available resources, ultimately improving the effectiveness of the guidance provided. Despite its importance, personalization in question answering remains underexplored, making it a timely and significant challenge for advancing interactive and user-centric information systems, with clear applications in web search and emerging systems such as AI mode in search. [^1] https://search.google/ways-to-search/ai-mode/
Figure 1. Pathways of Thoughts overview: It explores multiple directions of thinking about answering the question, generates a response for each, and mix them into a response.
Augmenting LLM inputs with personalized context of the user have proven effective for personalized text generation (Salemi et al., 2024b, a; Kumar et al., 2024; Li et al., 2023). This can be further enhanced by training the LLM to perform reasoning over the personalized context to extract the user’s preferences before response generation (Salemi et al., 2025). However, in most real-world cases, fine-tuning the LLM is often infeasible for two key reasons. On one hand, many LLMs are black boxes with frozen parameters, offering no control over their internal mechanisms. This is typical of popular applications like ChatGPT [^2] https://chat.openai.com/ and Gemini [^3] https://gemini.google.com/, which are only accessible through APIs or text interfaces. On the other hand, even with full access to model parameters, the fine-tuning process requires an inhibiting amount of computational resources for individual users. Consequently, a practical way to influence the model’s output is by applying advanced strategies during inference. Recognizing this, our paper focuses on this scenario and proposes an inference-stage strategy to achieve better personalization.
Achieving effective personalization presents its own unique challenges. First, personalized question answering fundamentally differs from other QA tasks in its definition of correctness. A response must be correct from the user’s specific perspective, considering their preferences, context, and background. Even when two answers are factually accurate, one may be far more suitable for a user’s specific information needs. Second, user preferences are often not explicitly stated and must be inferred from personalized context, such as profiles or prior interactions. This context is frequently lengthy and noisy, making it difficult for a model to extract the relevant signals. These issues highlight that simply augmenting an LLM’s input with user context is often insufficient. Effective personalization requires the model to perform complex reasoning over noisy, implicit user data to deduce user preferences before generating a response.
To address the aforementioned challenges of personalized question answering, this paper introduces Pathways of Thoughts (PoT), demonstrated in Figure 1. This approach, which can be applied to any LLM without additional training, enables the model to explore multiple thinking pathways when generating responses, each leading to a distinct response. We formalize each thinking pathway as a Markov Decision Process (MDP), where the LLM serves as both the agent and the environment. At each step, the LLM, acting as the agent, selects the most suitable action from a predefined set of cognitive operations, such as planning, reasoning, personalizing, and revising. The LLM then assumes the role of the environment, executing the chosen action and updating the state with newly acquired knowledge. This iterative process allows the model to refine its understanding of the user’s intent, extract key preferences from the user profile, and explore diverse strategies for answering the question. The process continues until the model determines that no further actions are needed, producing an initial set of responses. To generate the final personalized response, the LLM mixes outputs from different thinking pathways into a single response, aligning them with the user’s preferences extracted from their profile. Pathways of Thoughts addresses the challenges of LLMs in personalized question answering by allowing the model to iteratively think about the next step in responding to the question and how the lengthy user context can be effectively utilized. Additionally, it explores different diverse directions and paths, considering a range of possible solutions for answering the question. Finally, aggregation ensures that the output captures the strengths of all explored pathways, resulting in a more personalized and aligned response.
We conduct our experiments on the LaMP-QA benchmark (Salemi and Zamani, 2025a), designed for evaluating personalized question answering in three diverse domains: (1) Art & Entertainment, (2) Lifestyle & Personal Development, and (3) Society & Culture. Our results demonstrate that PoT achieves the best performance across all categories, with statistically significant improvements, up to $13.1\%$ relative improvement on average compared to the baselines. Additionally, we investigate the impact of pathway length, the number of pathways, and various methods for diversifying pathways and aggregating responses on performance. We show that PoT generalizes to any long-context LLM without requiring additional training. Finally, human evaluation shows that PoT better satisfies users’ information seeking needs in $66\%$ of cases, compared to $19\%$ for the baseline.
2. Related Work
Personalization of LLMs
Personalization plays a central role in search, recommendation, and text generation (Fowler et al., 2015; Xue et al., 2009; Naumov et al., 2019; Salemi et al., 2024b). To enable personalization in LLMs, Salemi et al. (2024b) introduced a retrieval-augmented generation framework alongside the LaMP benchmark for short-form text, which was later extended to long-form content generation with LongLaMP (Kumar et al., 2024) and to personalized QA with LaMP-QA (Salemi and Zamani, 2025a). Personalized assistants have also been studied in several recent works (Li et al., 2023; Mysore et al., 2023; Lu et al., 2024; Zhang et al., 2024a). Proposed approaches span a variety of techniques, including training retrieval models with user feedback (Salemi et al., 2024a), optimizing LLMs with personalized supervision (Jang et al., 2023), and generating personalized prompts (Li et al., 2024). Parameter-efficient fine-tuning has emerged as another direction (Tan et al., 2024), including its integration with retrieval-augmented methods (Salemi and Zamani, 2024). Beyond model adaptation, reasoning and self-training strategies have been shown to improve personalization in text generation (Salemi et al., 2025). More recently, natural language feedback has proven effective for adapting systems to individual users in personalized QA (Salemi and Zamani, 2025b). Nevertheless, strategies for improving personalized QA at test time—without incurring the cost of additional training—remain underexplored, which is the focus of this work.
Reasoners & Agents for Problem Solving
Reasoning enables models to solve problems with step-by-step processing, known as chain-of-thought (CoT), enhancing performance in complex tasks like mathematical problem-solving, logical reasoning, and commonsense understanding (Wei et al., 2022; Liu et al., 2023; Yin et al., 2024). Training LLMs to perform intermediate reasoning steps before generating responses has been effective, as seen in GPT-o1 (OpenAI, 2024) and DeepSeek-R1 (DeepSeek-AI, 2025). Additionally, agentic and multi-turn or multi-agent systems that interact with external tools, code (Liu et al., 2025; Xie and Zou, 2024; Fu et al., 2024), simulated environments (Hao et al., 2023; Wang et al., 2024; Park et al., 2024; Yao et al., 2022), or with other agents (Guo et al., 2024) have been explored for complex problem-solving. Markov Decision Processes (MDP) have shown promise in reasoning for games and mathematical problems (Hao et al., 2023; Wang et al., 2024). However, these methods have not been applied to free-form text generation, particularly personalization which requires user-specific responses rather than objectively correct answers. This study adapts these approaches for personalized text generation, combining them with graph search instead of traditional tree search in MDPs.
Scaling Inference Compute
Recent advances in LLM reasoning for logic and mathematics show that increasing compute resources during inference can significantly enhance performance (Snell et al., 2024; Chen et al., 2024). This allows LLMs to explore answer spaces more thoroughly, improving accuracy in logical reasoning, math problem-solving, and code generation (Bi et al., 2024; Zhang et al., 2024b; Brown et al., 2024). While previous research has focused on math, coding, and logic, its potential for free-form text generation, especially personalized question answering, is underexplored. Our work addresses this by using increased inference compute to better navigate response spaces, enhancing personalization.
3. Problem Formulation
We assume a question $x_{u}$ originates from a user $u$, who has a set of $n_{u}$ personalized information elements, denoted as $P_{u}=\{p_{i}\}_{i=1}^{n_{u}}$. Following previous work that utilizes long-term user history as the user profile (Salemi and Zamani, 2025a; Salemi et al., 2024b), $P_{u}$ consists of the user’s previously asked questions along with the their question narrative. The objective is to use this personalized information, along with the question $x_{u}$, to generate a personalized response for the user using an LLM $\pi$, formulated as $\hat{y}_{u}=\pi(x_{u},P_{u})$. To evaluate the personalized response generated for user $u$’s question, we assume access to a set of $n_{x_{u}}$ aspects $E_{x_{u}}=\{e_{i}\}_{i=1}^{n_{x_{u}}}$ that user $u$ expects to be addressed in the response, extracted from a question narrative $r_{x_{u}}$ written by the user $u$. Note that these aspects are used solely for evaluation and are not accessible to the model during response generation. Finally, a metric $\mu(x_{u},\hat{y}_{u},E_{x_{u}},r_{u})$ is used to score the generated response based on the extent to which these aspects are covered.
4. Pathways of Thoughts
To address the limitations of LLMs in personalized question answering, as discussed in Section 1, we introduce Pathways of Thoughts (PoT), as illustrated in Figure 1. In this approach, the LLM explores multiple thinking pathways to solve the problem, each leading to a separate response. We model the thinking process in each pathway as a Markov Decision Process (MDP), where the LLM functions as both the agent and the environment. At each step, the LLM, acting as the agent, selects an action from a set of fundamental cognitive actions, such as planning, reasoning, and revising. Then, the same LLM acts as the environment to execute the chosen action, updating the state with the newly acquired knowledge. This iterative process continues until no further actions are needed, leading to the generation of a final response. To produce a single personalized response based on all pathways, the LLM aggregates the generated responses from different thinking pathways, leveraging the strengths of each while aligning with the user’s preferences extracted from their profile. The details of this approach are provided in Algorithm 1. The following section elaborates on this method.
Algorithm 1 Pathways of Thoughts Implementation using MDP and Global Search.
1:prompt $x_{u}$, profile $P_{u}$, max path length $T$, number of pathways NN, backbone LLM $\pi$
2:personalized output $\hat{y}_{x_{u}}$}
[…]
4:for $i=0$ to NN do $\triangleright$ Generate NN pathways of thoughts
5:  $t=0$
6:  $s_{t}=\text{init\_prompt}(x_{u},P_{u})$ $\triangleright$ Prompt in Figure 11
7:  while $t<T$ do $\triangleright$ Thinking for maximum $T$ steps
8:   $A_{s_{t}}=\text{possible\_actions}(s_{t},t)$ $\triangleright$ Find the possible actions for state $s_{t}$.
9:   $\bar{a}_{p}=\text{action\_selection\_prompt}(s_{t},A_{s_{t}})$ $\triangleright$ Prompt in Figure 7
10:   $a_{t}=\pi(s_{t};\bar{a}_{p})$ $\triangleright$ Selecting the action (LLM as agent)
11:   $\hat{y}_{t}=\pi(s_{t};\bar{a}_{p};a_{t})$ $\triangleright$ Performing the action (LLM as environment)
12:   $s_{t+1}=s_{t};\bar{a}_{p};a_{t};\hat{y}_{t}$ $\triangleright$ new state by concatenating input/outputs in this time step
13:   if $a_{t}$ = finalize then
14:     $R_{\text{PoT}}=R_{\text{PoT}}\cup\{\hat{y}_{t}\}$ $\triangleright$ Capturing the final response of this pathway
15:     break
16:   end if
17:   $t=t+1$ $\triangleright$ updating the step variable
18:  end while
19:end for
[…]
24:return $\hat{y}_{x_{u}}$} $\triangleright$ Returning final response
4.1. Thinking as a Markov Decision Process
The thinking process in the human mind involves leveraging and integrating a set of fundamental cognitive skills—such as reasoning, planning, and revision—that collectively form the foundation of structured thought (Frensch and Funke, 2002; Sternberg and Ben-Zeev, 2001). Drawing inspiration from sequential decision making in the human mind, we model the thinking process of LLMs as an MDP. Given the current thoughts and steps taken to generate a response, the model selects the next fundamental action from a predefined set to enhance its understanding of the question, refine its reasoning, or improve the response.
An MDP is defined as a quadruple $(S,A_{s_{t}},P_{a_{t}},R_{a_{t}})$, where $S$ is the set of states, $A_{s_{t}}$ represents the set of available actions in state $s_{t}$, $P_{a_{t}}(s_{t+1}\mid s_{t})$ defines the probability of transitioning from state $s_{t}$ to $s_{t+1}$ given action $a_{t}$}, and $R_{a_{t}}(s_{t+1},s_{t})$ specifies the reward function, which assigns a reward for taking action $a_{t}$} in state $s_{t}$ and transitioning to $s_{t+1}$. In this setting, the agent $\pi$ aims to assign a probability $p_{\pi}({a_{t}}|s_{t})$ to each action based on the current state $s_{t}$.
To generate a personalized response to the question $x_{u}$ from user $u$, we assume the agent $\pi$ starts from an initial state $s_{0}$, which consists of the question $q_{u}$ and the user profile $P_{u}$. The agent iteratively selects the most probable action at each step using the prompt shown in Figure 7 in Appendix A, continuing this process until it either chooses a halting action (as determined by the LLM) or reaches the maximum number of allowed actions, denoted as $T$. Since our task involves text generation and lacks an external environment enforcing termination, the agent itself determines when to stop or continues until the predefined action limit is reached. Algorithm 1 (lines 3-17) demonstrates the algorithm. The following outlines the definitions of each MDP component in our framework:
Set of States ($S$):
The state of the system is defined as the concatenation of the inputs, outputs, and decisions of the model into a sequence of tokens using a predefined prompt format. This representation captures the full context of the system like a conversation. The initial state of the system consists of the user’s question $x_{u}$, the personalized context $P_{u}$ about user $u$, and a set of instructions specifying the actions and the model’s expected behavior for each action, as shown in Figure 11 in Appendix A. Given this setup, the set of all possible states corresponds to the set of all textual sequences that this initial prompt is their prefix. As a result, the state space forms an uncountable discrete set.
Set of Possible Actions for State $s_{t}$ ($A_{s_{t}}$):
We consider the following fundamental actions that the agent can choose from in thinking process: answering, planning, checking for personalization, personalizing, reasoning, clarifying, summarizing, revising, and finalizing. [^4] This action, finalizing, also referred to as halting, terminates the process. The detailed definition and output format of each action is provided in Appendix A.1. In the initial state, the agent is restricted to selecting the planning action, as it must first plan the steps required to address the user’s question. In subsequent states, the agent can choose from all the actions; however, the model can only select the revising action if it has previously selected the answering action, as revision typically follows the generation of a response.
Transitioning from State $s_{t}$ to $s_{t+1}$ with Action $a_{t}$} ($P_{a_{t}}(s_{t+1}\mid s_{t})$):
We use the LLM as both the agent and the environment simultaneously: when it selects an action, it acts as the agent, and when it executes the action, it acts as the environment. This dual role allows the LLM to control the flow of thinking and performing it simultaneously. Accordingly, with the agent decision to perform action $a_{t}$ at time step $t$, the LLM receives the state $s_{t}$ and action $a_{t}$ as input and generates the output $\hat{y}_{t}$. The concatenation of $s_{t}$, $a_{t}$, and $\hat{y}_{t}$ forms the next state $s_{t+1}$. Thus, the transition probability $P_{a_{t}}(s_{t+1}\mid s_{t})$ is defined by the LLM itself.
Reward Function ($R_{a}(s_{t+1},s_{t})$)
Since our agent is used without training due to optimization costs, an explicit reward function is not required; we assume the agent makes reasonable choices. However, in general, an outcome reward function can be employed to assign rewards to actions based on the evaluation of the LLM’s final response to the question.
4.2. Multi-Directional Thinking by Diverse Planning and Mixture-of-Responses
Table 1. Performance on LaMP-QA. $\dagger$ and $\ddagger$ show statistically significant improvement over the single and multiple inference best baseline, respectively, using t-test ($p<0.05$). In this experiment, Gemini 1.5 Pro has been used as the backbone LLM for all the methods.
Method Arts & Lifestyle & Personal Society & Average Entertainment Development Culture (macro) Single Inference (1) No Personalization 0.1741 0.2863 0.2996 0.2533 (2) In-Context Personalization 0.1722 0.3022 0.3482 0.2742 (3) In-Context Personalization w/ CoT 0.2119 0.3862 0.4230 0.3403 (4) Pathways of Thoughts ($N=1$, $T=8$) 0.2499 0.3966 0.4490 0.3651 Multiple Inference (5) In-Context Personalization w/ Best-of-32 0.2248 0.3753 0.4303 0.3434 (6) In-Context Personalization w/ CoT & Best-of-32 0.2607 0.4380 0.4789 0.3925 (7) Tree of Thoughts ($K=32$, $B=2$) 0.2054 0.4178 0.4546 0.3592 (8) Pathways of Thoughts ($N=16$, $T=8$) 0.2999†‡ 0.4960†‡ 0.5362†‡ 0.4440†‡
We observed that when the agent selects planning as the first action, the subsequent steps consistently follow the outline of the generated plan. This suggests that encouraging multi-directional thinking about responding to the question can be achieved by generating a variety of plans, each exploring a different strategy for addressing the question. We study two methods for diverse planning:
  • Initial State Alteration: If the initial state is modified before executing the planning action, the LLM can produce different plans, as the changes in personalized context influences the planning process. This can be modeled by varying the amount of past user history accessible to the agent, effectively addressing the challenge of long personal context. To implement this, we randomly select $N$ subsets of the user profile $P_{u}$, each of size $\tau\times|P_{u}|$, where $0\leq\tau\leq 1$. Starting from these $N$ diverse initial states lead to different planning trajectories and thinking pathways.
  • Planning Action Variation: Another way to generate different plans from the same initial state is to use a higher sampling temperature $0\leq\tau\leq 1$ during plan generation, which increases the randomness in the generated plans, leading to more diverse outputs. Thus, we generate $N$ different plans from the same initial state (created using the whole profile $P_{u}$) by setting a high temperature $\tau$, leading to different thinking pathways.
Each thinking pathway results in a response, and together they form a set of response proposals $R_{\text{PoT}}=\{r_{i}\}_{i=1}^{N}$. To produce the final response $\hat{y}_{x_{u}}$ for the user, we need to ensure that it is the most suitable one for the user. To achieve this, we first use the LLM $\pi$ to extract the important aspects $I_{u}$ for user $u$ from their user profile $P_{u}$ using the prompt shown in Figure 8 in Appendix A.2. This serves as an intermediate step to learn about the user’s individual preferences. Based on this, there are two methods to generate the final response:
  • Best-of-N: Given the extracted important aspects $I_{u}$ for the user $u$ and the input $x_{u}$, the LLM $\pi$ selects the best response from the set of responses resulted from multi-directional thinking $R_{\text{PoT}}$, using the prompt shown in Figure 9 in Appendix A.2.
  • Mixture-of-N: Given the extracted important aspects $I_{u}$ for the user $u$ and the input $x_{u}$, the LLM $\pi$ combines different aspects of the generated responses from the set of multi-directional thinking outputs $R_{\text{PoT}}$. The final response is denoted as $\hat{y}_{x_{u}}=\pi(x_{u},R_{\text{PoT}},I_{u})$, using the prompt shown in Figure 10 in Appendix A.2. This mimics how humans solve problems by exploring multiple solutions and synthesizing a new solution that incorporates the strengths of each solution. This is particularly beneficial for open-ended text generation, where there is no definitive right or wrong response. By aggregating the strengths of all responses, it generates a response that aligns with user preferences, ensuring a more personalized output. In PoT, we use this approach as the primary response aggregation method.
Figure 2. Effect of the maximum number of actions per pathway on the performance.
Figure 3. Effect of the number of pathways on the performance.
5. Experiments
5.1. Experimental Setup
Datasets
We conduct experiments on the LaMP-QA benchmark (Salemi and Zamani, 2025a), the only publicly available dataset for personalized QA. LaMP-QA covers three diverse domains: (1) Art & Entertainment (767 questions), (2) Lifestyle & Personal Development (989 questions), and (3) Society & Culture (1074 questions). [^5] Other datasets for personalized text generation, such as LaMP (Salemi et al., 2024b) and LongLaMP (Kumar et al., 2024), are not suitable for our setting for two main reasons. First, they primarily target open-ended content generation, which differs from the QA focus of this work (Section 1). Second, unlike LaMP-QA, which provides explicit ground-truth annotations of user expectations, they include only a single reference response. This restricts their ability to capture the diversity of valid personalized outputs and can lead to misalignment with user expectations, as noted by Salemi and Zamani (2025a). These limitations are particularly problematic for our feedback-based approach, where convergence to a single reference response is both infeasible and suboptimal for reliable evaluation. Each LaMP-QA example contains a user query, the user’s question history (serving as a profile), a narrative that reflects the user’s perspective and intent, and a set of personalized rubrics specifying the aspects an ideal response should address. For evaluation, we adopt the procedure of Salemi and Zamani (2025a), but replace their suggested evaluator (Qwen-2.5, 32B) with Gemini 1.5 Pro, a more capable model. For each question, the evaluator checks whether the generated response addresses each personalized aspect, assigning scores in the range $[0,2]$. These are normalized to $[0,1]$, and the final response score is computed as the mean normalized score across all aspects. For details, we refer the reader to Salemi and Zamani (2025a). In one experiment, we corroborate our findings through a side-by-side human evaluation comparing PoT to the strongest baseline.
PoT’s Configurations
We use Gemini 1.5 Pro as the backbone LLM. To show the generalizability of PoT, we use GPT-4o-mini [^6] Available at: https://platform.openai.com/docs/models/gpt-4o-mini (OpenAI, 2024) using our most representative experiment configurations. For single pathways, we use Nucleus sampling (Holtzman et al., 2020) with a temperature of $\tau=0.1$. For multi-directional pathways, we use the same setting, except for the Planning Action Variation, where a temperature of $\tau=0.9$ is used exclusively for the planning action. For generating diverse plans, we use Planning Action Variation by default and Mixture-of-N for response aggregation, unless otherwise specified. We set the context length to $32k$, the maximum number of actions per pathway to $T=8$ , and the number of pathways to $N=16$. In ablations, we reduce the number of pathways to $N=4$ to save costs, unless otherwise specified. For personalized context, we use the 10 recent questions of the user with the same order as they appear in the LaMP-QA (Salemi and Zamani, 2025a) dataset.
Baselines
We use 6 baselines: (1) LLM without personalized context to directly answers the question, (2) LLM with access to personalized context, and (3) LLM with access to personalized context that first performs reasoning over the context to extract important aspects relevant to the question before generating the response. The two latter use the prompts shown in Figures 12 and 13 in Appendix B. These baselines all use Nucleus sampling with a temperature of $\tau=0.1$. We extend the latter two baselines by doing multiple inferences and Best-of-N (Ichihara et al., 2025; Brown et al., 2024) to evaluate their performance when test-time compute scales, to form baselines (4) and (5). Additionally, we include the (6) Tree of Thoughts (ToT) (Yao et al., 2023) as a baseline, leveraging its test-time compute scaling. We follow their setup, particularly in the creative writing task, as it is the closest task they studied to personalized question answering. This setup consists of an intermediate planning step, where a set of plans is generated and the best plan is selected, followed by a final response generation step, where a set of responses is generated and the best response is chosen. For planning, we use the prompts shown in Figure 14, and for response generation, we refer to Figure 15, both in Appendix B. We observed that the baselines generate, on average, half the number of tokens compared to PoT{PoT}. To ensure a fair comparison, we set the best of N baselines to generate double the number of responses, with $N=32$ and a sampling temperature of $\tau=0.9$ to ensure diverse responses. For the Best-of-N approach, we utilize the method introduced in Section 4.2. For ToT, following Yao et al. (2023) and their setup for the creative writing task that is the most similar to our task, we set the tree depth to $B=2$. For fair comparison with PoT, we generate $N=32$ plans and responses at each node with a sampling temperature of $\tau=0.9$. The prompts for selecting the best plan is shown in Figure 16 in Appendix B. For selecting the best response, we utilize the method introduced in Section 4.2.
5.2. Empirical Findings
Comparison with baselines
The results of this experiment are reported in Table 1. The findings show that PoT statistically significantly outperforms all the baselines across all categories, achieving a 13.1% relative improvement on average over the best baseline (row 6). This clearly demonstrates the superiority of PoT compared to the baselines. To compare our thinking-as-an-MDP method with baselines, we also report the results using only a single inference (rows 1-4). The results from this experiment show that our method (row 4) outperforms all single inference baselines (rows 1-3), with statistically significant improvements in two out of three categories and overall performance. Specifically, our method achieves a 7.2% relative improvement in average performance, suggesting that our thinking approach is significantly more effective than CoT.
Effect of the maximum number of actions in PoT
For this, we focus on a single inference ($N=1$) using PoT and vary the maximum number of actions it can perform. The performance across all categories is shown in Figure 2. The findings show that increasing the number of actions improves performance up to a certain point (8 actions per pathway), after which performance declines. This occurs because, with too few actions, the LLM cannot construct effective pathways to solve the problem. Conversely, when given too many actions, the LLM may overthink the problem, performing unnecessary actions that degrade performance.
Figure 4. Effect of planning diversification and response aggregation methods on the performance.
Figure 5. Performance of Pathways of Thoughts and baselines with GPT-4o-mini as the backbone LLM.
Effect of the number of pathways in PoT
We set the maximum number of actions per pathway to $T=8$ and vary the number of pathways. The performance across all categories is shown in Figure 3. The findings indicate that increasing the number of pathways consistently improves performance; however, as the number grows, the improvements become smaller. This occurs because there is a limited set of different ways for answering the question, and after a certain point, the solutions start to repeat, leading to diminishing returns in performance with additional pathways.
Effect of different plan diversification and response aggregation in PoT
We evaluate all possible combinations of the methods introduced in Section 4.2. Specifically, we consider four configurations: 1) Planning Action Variation w/ Best-of-N, 2) Planning Action Variation w/ Mixture-of-N, 3) Initial State Alteration w/ Best-of-N, and 4) Initial State Alteration w/ Mixture-of-N. Additionally, we explore different values of the hyperparameter $\tau$ for both Planning Action Variation and Initial State Alteration. The results of this experiment are shown in Figure 4. Our analysis reveals several key observations. Most notably, Mixture-of-N consistently outperforms Best-of-N regardless of the plan diversification method. This can be attributed to the fact that Best-of-N selects only one response from the generated candidates, whereas Mixture-of-N enables the combination of strengths from multiple responses, resulting in a more comprehensive and higher-quality output. we observe that increasing the hyperparameter $\tau$ improves performance. In Planning Action Variation, a higher $\tau$ leads to a more diverse set of generated plans due to higher sampling temperature, resulting in more varied thinking pathways that may produce more tailored responses. In Initial State Alteration, increasing $\tau$ provides each pathway with more user-specific information, enabling more personalized responses. Comparing the two, Planning Action Variation generally outperforms Initial State Alteration in terms of average performance. However, as $\tau$ increases, the performance gap narrows. Overall, Planning Action Variation w/ Mixture-of-N achieves the best performance for personalized question answering.
Figure 6. A case study of Pathways of Thoughts with $N=16$ and $T=8$ for question: What’s the best way for me to bring up a topic of political concern if the business I work for is minority-led?
PoT’s generalizability and independence of LLM
In the same setting as Gemini 1.5 pro, we apply PoT and the best performing baselines to another backbone LLM, GPT-4o-mini, to demonstrate its effectiveness across different LLMs with long context windows. The average performance across all categories is presented in Figure 5. The results indicate that, similar to Gemini 1.5 Pro, GPT-4o-mini with PoT outperforms all baselines on all categories. It achieves a statistically significant 9.4% relative improvement on average across all datasets compared to the best baseline. These findings suggest that PoT enhances the performance of LLMs in personalized QA regardless of the underlying backbone LLM.
Human evaluation.
We randomly sample 100 examples from LaMP-QA and generate outputs using PoT and In-Context Personalization w/ CoT & Best-of-32. Human annotators evaluate the outputs based on their alignment with the the user’s question narrative. Each example is annotated twice, yielding an inter-annotator agreement Cohen’s kappa of $0.613$. The results indicate that annotators preferred PoT in 66% of the cases, the baseline in 19%, and reported tie in 15% of cases. These results show that PoT achieves considerably higher alignment with human preferences.
5.3. Case Study
We observe that PoT ($N=16$, $T=8$) generates $1943$ unique pathways, defined by distinct action sequences chosen by the LLM. Specifically, Art & Entertainment yields $1109$, Lifestyle & Personal Development produces $814$, and Society & Culture results in $1182$ unique pathways. On average, each example generates $11.1$ unique pathways with on average $7.19$ actions. As a case study demonstrating the effectiveness of PoT, we present an example in Figure 6:
  • Thinking Pathways & Generated Responses: To conserve space, we present in Figure 6 only the sequence of steps taken in each pathway, rather than the full generated text, and the response in each pathway. Each pathway corresponds to a distinct reasoning trajectory generated by PoT. For example, in the illustrated case, PoT selects different pathways depending on whether personalization is required. Some pathways generate a direct response without personalization (e.g., Pathway 7), while others involve multiple rounds of reasoning that integrate various aspects of the query with the user profile before producing an answer (Pathways 1–6, 9–11, 13–15). Certain pathways also involve revision (Pathways 12, 15, and 16) or clarification (Pathway 4). Notably, we observed that for some trajectories PoT omit personalization entirely (Paths 7, 8, and 11). This shows that PoT deliberately explores a diverse range of reasoning strategies, producing responses from multiple perspectives that are subsequently combined into a final answer encompassing the advantages of each. Although responses across pathways share common elements, each retains distinctive characteristics, reflecting the variation in underlying reasoning strategies.
  • Final Response: Figure 6 shows the final response to the user’s question, produced by aggregating and integrating components from multiple pathway outputs into a single coherent answer. The contributions of different pathways are indicated in brackets ([]). This shows how PoT leverages the complementary strengths of diverse reasonings, combining them into a unified response that aligns with the user’s preferences and key information needs.
6. Conclusions
This paper addresses the critical yet under-explored problem of personalized question answering. we propose Pathways of Thoughts (PoT), a novel approach that models the LLM’s thinking process as a Markov Decision Process (MDP) during test time. By enabling the exploration of multiple thinking pathways and aggregating diverse responses, PoT effectively captures user preferences to generate highly personalized answers. Our experimental results on LaMP-QA demonstrate that PoT achieves substantial improvements over baseline methods, with performance gains of up to 13.1%. Furthermore, human evaluation confirms the effectiveness of our approach, showing a strong user preference over existing baselines. This paper focuses solely on personalized question answering; however Pathways of Thoughts is a general framework and future work can explore the potential of Pathways of Thoughts for long-form text generation tasks beyond personalization.
Appendix A PoT’s Implementation Details
A.1. Actions Definition
We define an action as a means by which the agent interacts with the text generation environment, specifically the LLM, by instructing it to perform a particular operation. Each action is associated with a specific output format that the LLM generates in response to the instruction. In this paper, we consider the following actions as the operations the model can take to respond to a question:
  • answering: This action instructs the model to generate a personalized response to the question, provided that personalization can enhance the response. It should be selected when the model is prepared to answer the question, ensuring that all necessary information is available and there is no ambiguity.
  • planning: This action instructs the model to generate a plan of steps required to address all aspects of the question. It should be executed when the question has multiple components, and the model recognizes that planning the next steps is necessary to adequately respond to the user’s query.
  • checking for personalization: This action asks the model to evaluate whether the user’s question can benefit from personalization. It is useful when determining whether personalizing the response is necessary or if a generic response would suffice.
  • personalizing: This action asks the model to identify relevant information from the user profile and summarize how it can be used to answer the question. It is executed when personalized information is necessary to enhance the quality of the response.
  • reasoning: This action prompts the model to break down one aspect of the question and think through it step by step. It is executed when questions involve multiple aspects, and the model needs to focus on and consider how to address a specific aspect.
  • clarifying: This action prompts the model to assess whether the question is ambiguous and determine how it can be clarified. It is executed when the model identifies that the current information is insufficient or unclear, requiring further elaboration.
  • summarizing: This action instructs the model to summarize all available information and findings. It should be executed when the model determines that the current information needs to be consolidated or synthesized to ensure a clearer understanding before continuing with the response generation.
  • revising: This action instructs the model to revise the current response to the question. It should be executed when the model has already generated an initial response but believes the response can be improved or adjusted to better address the user’s query or to incorporate additional insights or information.
  • finalizing: This action finalizes the response. It should be executed when the model believes the response is complete, accurate, and addresses the user’s query. No further actions or revisions are necessary, and the response is ready to be presented.
A.2. Details of Response Aggregation Methods
Figure 12. Prompt for the In-Context Personalization.
Figure 13. Prompt for In-Context Personalization with CoT.
Figure 14. Prompt for the planning step in ToT.
An essential component of Pathways of Thoughts involves aggregating the candidate responses generated across different thinking pathways into a single output. Each pathway produces a candidate response, collectively forming a set of response proposals, denoted as $R_{\text{PoT}}$. To generate the final response $\hat{y}_{x_{u}}$ personalized for the user, it is crucial to select the most suitable candidate. To achieve this, we first employ the LLM $\pi$ to extract the key aspects $I_{u}$ relevant to user $u$ from their user profile $P_{u}$. This extraction process is guided by the prompt illustrated in Figure 8, acting as an intermediate step to capture the user’s individual preferences. Based on this, there are two methods to generate the final response:
  • Best-of-N: Given the extracted important aspects for the user and the input, the LLM selects the optimal response from the set of candidate responses generated through multi-directional thinking. This is guided by the prompt shown in Figure 9.
  • Mixture-of-N: Given the extracted important aspects for the user and the input, the LLM integrates different aspects of the candidate responses from the set of multi-directional thinking outputs, guided by the prompt in Figure 10. This method emulates human problem-solving by exploring multiple solutions and synthesizing a new response that leverages the strengths of all candidates.
Figure 15. Prompt for the response generation step in ToT.
Figure 16. Prompt for the plan selection step in Tree of Thoughts (Yao et al., 2023). The important aspects extracted for the user are generated using the prompt shown in Figure 8.
Appendix B Baselines Prompts
Figures 12 and 13 present the prompts used for the personalized baselines without and with chain-of-thought reasoning, respectively. For the Tree of Thoughts baseline, Figure 14 shows the planning prompt, Figure 15 illustrates the response generation prompt, and Figure 16 provides the plan selection prompt.